OpenStudy (anonymous):

omg can someone solve this (with steps shown) (x+6)(x-7)(x-8)(x+9)=0 by factoring?

OpenStudy (anonymous):

that is factored

OpenStudy (anonymous):

Already Factored! x+6 = 0 x-7 = 0 x-8 = 0 x+9 = 0

OpenStudy (anonymous):

factored already

OpenStudy (anonymous):

x = -6, 7, 8, -9

myininaya (myininaya):

the solutions of \[(x-a)(x+b)(x-c)(x+d)=0\] are \[x=a,-b,c,-d\]